I'm trying to the determine the base of a right hand triangle that has a base of 4 and height 4 - I was told the answer is b-5.66 where does the 5.66 come from?
Answer by nerdybill(7384)

They are applying the Pythagorean theorem:
a^2 + b^2 = c^2
where
'a' and 'b' are the legs of a right triangle
and 'c' is the hypotenuse (longest side)
4^2 + 4^2 = c^2
16 + 16 = c^2
32 = c^2
sqrt(32) = c
5.66 = c